OpenStudy (anonymous):

Solve the equation step by step, list each step, along with the property you used during that step, -5×1×11×4 I will be very thankfull to anyone who can answer this for me plus I'll give a medal :-)

OpenStudy (anonymous):

There is no \(=\)? This is one messed up "equation"

OpenStudy (anonymous):

That's how it shows up in the worksheet B-)

OpenStudy (anonymous):

multiply them together. if you want to takes steps it would be \[-5\times 1\times 11\times 4=-5\times 11\times 4=-55\times 4=-220\]

OpenStudy (anonymous):

but as @wio said, this is not an equation, it is a computation. an equation needs an equal sign